Suppose $200 worth of government bonds are sold on the market to raise public funds, and that $150 of this is diverted from current private consumption spending, and $50 is from current private investment spending. The tax rate on the returns to private investment is 20%.

a) calculate

1. the total cost to the economy of raising the $200 of public funds

2. the deadweight loss

3. the marginal cost per dollar of public funds raised.

b) if this project is to be funded entirely from public funds and the present value of its benefits and costs are $210 and $1200 respectively, show how you would apply the NPV decision rule in this case, factoring in the cost of raising the $200 of public funds?
Answer #1

Answer Given data. $200 worth o government bonds me sold on the market to raie public fondo. $150 of this as diverted from cuThe dead weight loss = Total cost to the Economy - Government bond = 21215 - 200 $12.5. 3 of public The marginal cost per dol@ (Cost + dead blight Cost : ght (08) =(C+DYC = 200 +1215 200 = 21215/200 = 1.06. this we obsence Blc < (C+D)/e. from that So